What will be the oxidation number of Ni in.  K2[Ni(CN)4]?

Grade:11

3 Answers

Arun
5 years ago

In the compound K2(Ni(CN)4), the anion has a charge of -2, because the compounds contains 2 K+ ions, and thus the charge of the anion should balance out the charge of the cations.

We know that CN has a charge of -1. Therefore we can find the oxidation number of Ni through simple calculation.

If we allow the charge of Ni to be x, then:

x + 4(-1)= -2

x = +2

Therefore the oxidation number of Ni in K2(Ni(CN)4) is +2.

bhavana
5 years ago
It is +2 
Since the over all charge on complex part is -2 
Therefore -2=x+(-1*4)
So x (charge of Ni is +2
na batau
5 years ago
 

In the compound K2(Ni(CN)4), the anion has a charge of -2, because the compounds contains 2 K+ ions, and thus the charge of the anion should balance out the charge of the cations.

We know that CN has a charge of -1. Therefore we can find the oxidation number of Ni through simple calculation.

If we allow the charge of Ni to be x, then:

x + 4(-1)= -2

x = +2

Therefore the oxidation number of Ni in K2(Ni(CN)4) is +2.